2. A random variable X can have values -4. - 1, 2, 3 and 4, each with probability 1/5 Find: the density function f_{Y} . the distribution function F_{y} , the mean E(Y) . and the variance sigma_{y} ^ 2 , of the random Y = 3X ^ 3

1
Expert's answer
2022-04-06T13:57:00-0400

"P\\left( Y=-192 \\right) =P\\left( X=-4 \\right) =\\frac{1}{5}\\\\P\\left( Y=-3 \\right) =P\\left( X=-1 \\right) =\\frac{1}{5}\\\\P\\left( Y=24 \\right) =P\\left( X=2 \\right) =\\frac{1}{5}\\\\P\\left( Y=81 \\right) =P\\left( X=3 \\right) =\\frac{1}{5}\\\\P\\left( Y=192 \\right) =P\\left( X=4 \\right) =\\frac{1}{5}\\\\f_Y\\left( y \\right) =\\left\\{ \\begin{array}{c}\t0.2,y\\in \\left\\{ -192,-3,24,81,192 \\right\\}\\\\\t0,otherwise\\\\\\end{array} \\right. \\\\F_Y\\left( y \\right) =\\left\\{ \\begin{array}{c}\t0,y<-192\\\\\t0.2,-192\\leqslant y<-3\\\\\t0.4,-3\\leqslant y<24\\\\\t0.6,24\\leqslant y<81\\\\\t0.8,81\\leqslant y<192\\\\\t1,y\\geqslant 192\\\\\\end{array} \\right. \\\\EY=\\frac{-192-3+24+81+192}{5}=20.4\\\\EY^2=\\frac{\\left( -192-20.4 \\right) ^2+\\left( -3-20.4 \\right) ^2+\\left( 24-20.4 \\right) ^2+\\left( 81-20.4 \\right) ^2+\\left( 192-20.4 \\right) ^2}{5}=15758.6\\\\var\\left( Y \\right) =15758.6-20.4^2=15342.5"